THE HEIGHT OF A TRIANGLE IS 3 METERS GREATER THAN THE BASE. THE AREA OF A TRIANGLE IS 35 SQUARE METERS. FIND THE LENGTH OF THE BASE AND THE HEIGHT OF THE TRIANGLE.
Answer by Cromlix(4381)

Represent the height as (x + 3) and the base as x.
Area of triangle = 1/2 (base x height) = 35 m^2
1/2 (x times (x + 3)) = x^2 + 3/2.
x^2 + 3/2 = 35 m^2
x^2 + 3x = 35 x 2
x^2 + 3x = 70
x^2 + 3x - 70 = 0
(x + 10)(x - 7)
The answer is x = 7
Height = 10m and base = 7m